Assume a linear demand function of the form:

Qd = 120 - 5P and a linear supply curve of the form:

Qs = -30 + 10P

Using these demand and supply functions, answer the following questions.


1
Expert's answer
2022-04-17T17:03:50-0400

"Q_d=120-5P,"

"Q_s=-30+10P,"

"Q_d=Q_s,"

"120-5P=-30+10P,"

"150=15P,"

"P=10,"

"Q_d=Q_s=70."